7. Given Demand fn. Q = 100 – 0.2P (Or P = 500 – 5Q), Cost fn. TC = 50 + 20Q + Q2

A. Find the profit maximizing price and output of a perfectly competitive market,
B. Assuming, a firm is monopoly
C. Compare the results
1
Expert's answer
2021-09-08T14:14:46-0400

A. The condition for profit maximization for a competing firm is to choose such a volume of production that the price is equal to the marginal cost

"P=MC \\

500-5Q = MC=TC’=20+2Q \\

500-5Q=20+2Q \\

480=6Q \\

Q= \frac{480}{6}=80 \\

P\; (price) = 500-5\times 80 = 100 \\

P \;(profit \; max) = TR-TC \\

TR=P \times Q \\

P(profit \; max) = 100 \times 80 - (50+20 \times 80 + 80^2) =8000 -8050 =-50"


B. The condition for maximizing profits in the monopoly market is the equality of marginal costs and marginal revenue:

"MC=MR \\

MC=TC’ = 20 +2Q \\

TR=Q(500-5Q) = 500Q -5Q^2 \\

MR = TR’ = 500 – 10Q \\

20+2Q=500 -10Q \\

12Q=480 \\

Q=\frac{480}{12}=40 \\

P(prise) = 500 -5 \times 40 = 300 \\

P(profit \; max) = TR-TC = P \times Q - (50 + 20Q + Q^2) \\

= 300 \times 40 50 20 \times 40-40^2 \\

= 12000 -50- 800 – 1600 \\

= 9550"


C. The firm will incur losses when releasing goods in a competitive market. Effective work is possible only in conditions of monopoly.